A person stands 6.00 m from a speaker, and 8.00 m from an identical speaker. What is the frequency of the second (n=2) interference minimum (destructive)?

• I have asked this question before and for 428.75 Hz as an answer, which is telling me is an incorrect answer, I am unaware how to get the actual answer

A person stands 6.00 m from a speaker, and 8.00 m from an identical speaker. What is the frequency of the second (n 2) interference minimum (destructive)? (Hint: v Af 343 m/s) (Unit - Hz) nter
Answer #1

For destructive interference, the expression is as follows -

|L1-L2|= (n+1/2)λ

As mentioned in the problem, L1 = 8.0 m, L2 = 6.0 m

For second interference to be minimum -

Put, n=1

So, we have -

(1+1/2)λ= |8.0 - 6.0| m = 2.0 m

=> λ= (2 x 2) / 3 = 1.33 m

Therefore, the value of the frequency, f=v/λ

= 343.0 / 1.33 = 258.0 Hz (Answer)